OpenStudy (anonymous):

Use the binomial expression (p + q)n to calculate a binomial distribution with n = 5 and p = 0.3.

OpenStudy (anonymous):

@thomaster ? @MathLegend ? @mathbrz ? @math_genius123 ?

OpenStudy (anonymous):

(p+q)n =(0.3+q)5 =1.5+5q distribute the 5 into what is in the ( ) 5*0.3=1.5 5*q=5q =1.5+5q

OpenStudy (anonymous):

Thank you so much!!!!!

OpenStudy (anonymous):

no problem I hope that was easy to understand! :)

OpenStudy (anonymous):

Definitely! :)